# MambaFlow: A Mamba-Centric Architecture for End-to-End Optical Flow Estimation

### v2

@@ Abstract
Recently, the Mamba architecture has demonstrated significant successes in various computer vision tasks, such as classification and segmentation. However, its application to optical flow estimation remains unexplored. In this paper, we introduce MambaFlow, a novel framework designed to leverage the high accuracy and efficiency of the Mamba architecture for capturing locally correlated features while preserving global information in end-to-end optical flow estimation. To our knowledge, MambaFlow is the first architecture centered around the Mamba design tailored specifically for optical flow estimation. It comprises two key components: (1) PolyMamba, which optimizes feature representation; and (2) PulseMamba, which facilitates efficient flow information dissemination. Our extensive experiments demonstrate that MambaFlow achieves remarkable results. On the Sintel benchmark, MambaFlow records an endpoint error (EPE) of 1.43 and an inference speed of 0.113 seconds, surpassing the state-of-the-art methods including GMFlow (with 18.9% lower EPE and 18.1% faster inference), SeparableFlow (5% lower EPE and 50.5% faster), CRAFT (1.11% lower EPE and 76.5% faster), and DIP (0.7% lower EPE and 77.2% faster)—demonstrating stronger potential for real-world deployment on resource-constrained devices. The source code will be made publicly available upon acceptance of the paper.
1 Introduction
@@
Optical flow estimation, a fundamental problem in computer vision, aims to compute the motion vectors of pixels between consecutive frames in a video sequence. It holds significant and widespread applications in areas such as action recognition [65, 56], video interpolation [16, 51, 28], autonomous driving [59], etc.
Architectures based on Convolutional Neural Networks (CNNs) have achieved pioneering progress in the field of optical flow. FlowNet [5], as the first end-to-end CNNs-based flow network, revealed the immense potential of this learning paradigm. However, it faced limitations in handling large displacements and complex motions. Subsequently, PWC-Net [44] improved estimation accuracy in complex motion scenarios by incorporating pyramid feature extraction and a cost-volume construction, which effectively captured multi-scale local features. Nevertheless, it still encountered challenges in large displacement scenarios. To address these issues, RAFT [45] introduced a recurrent iterative refinement mechanism, which significantly enhanced the accuracy of optical flow estimation. However, such innovation brought a new challenge, which its optimization process relies on numerous iterations, leading to increased parameter counts and longer inference time. Many subsequent variants of the RAFT algorithm have attempted upgrates, such as MS-RAFT [19] and SEA-RAFT [49], but their ability to handle large displacements and complex motions still requires further improvement. To tackle the aforementioned issues, the Transformer architecture, known for its powerful global modeling capabilities, has been introduced into optical flow estimation tasks [15, 42, 30, 43]. Most of them leverage attention mechanism to enhance feature representation, which capture long-range dependencies to significantly improve the accuracy. Notably, unlike optical flow algorithms based on RNNs for flow decoding [45, 20], GMFlow [54] utilizes the attention mechanism to iteratively decode the final optical flow, resulting in faster flow refinement. However, they encounter the quadratic computational complexity of Transformers both during the training and inference stages [54, 43, 15, 42, 30], which severely limits their efficiency in practical applications.
@@
Figure 1: End-point-error on Sintel (clean) vs. inference time (ms) and model params (M). All models were trained on FlyingChairs (C), FlyingThings3D (T), Sintel (S), KITTI (K), and HD1K (H) datasets, with testing performed using one NVIDIA GeForce RTX 3090 GPU. Compared to state-of-the-art models, our proposed model achieves comparable accuracy while significantly reducing inference time, demonstrating an optimal balance between computational efficiency and performance.
@@
Recently, Mamba [8, 10] has been widely used in various computer vision tasks, such as image classification [7, 57, 58],image segmentation [27, 67, 6], object detection [33, 32, 46], image generation [13, 23], video understanding [21, 24, 52], etc. Those remarkable achievements have demonstrated that Mamba can effectively address the challenges posed by CNNs and Transformers, including the local perception limitations of CNNs and the quadratic computational complexity of Transformers [55, 63]. It benefits from a selective state space mechanism, which can efficiently capture global information and handle complex spatiotemporal dependencies with linear complexity [63]. Specifically, Mamba has been conceptually summarized as being ideally suited for tasks involving long sequences and autoregression [60] due to its inherent RNN mechanism within the SSM [10, 9, 8].
@@
Inspired by Mamba’s capabilities for long sequence modeling and autoregression with linear complexity, we propose a novel optical flow estimation framework, MambaFlow, designed to address the challenges in CNNs or Transformer-based methods. As demonstrated in Fig. 1, our method significantly reduces inference time while maintaining remarkblely high accuracy.The framework introduces two core components: (1) the PolyMamba module, which utilizes Self-Mamba, Cross-Mamba, and Multi-Layer Perceptron (MLP) to globally optimize feature representations with lightweight computation, and (2) the PulseMamba module, which effectively combines motion features with Mamba’s autoregressive ability to iteratively generate flow information, addressing challenges such as large displacement and occluded regions.
@@
Specifically, we begins with a CNN backbone to extract robust features from the input image pairs similar to traditional practice [49, 54, 20]. Subsequently, we leverage the suggested PolyMamba to enhance the extracted features, allowing for the capture of complex global information and long-range dependencies. Building on the enhanced features, we further utilize a global matching mechanism [54, 64] to compute similarities for initial flow. At last, we employ PulseMamba to recurrently refine and yield the final precise flow [54]. Our algorithm uses Mamba as the core components in both the encoding and decoding stage, making it an end-to-end optical flow estimation algorithm centered around Mamba. By leveraging Mamba’s capabilities in global modeling and autoregressive properties, as well as its computationally efficient characteristics, the proposed approach achieves a good trade-off between accuracy and performance.
@@
  • We design the PulseMamba, which includes a motion encoder and a Mamba to facilitate flow propagation, addressing the computational overload, large displacement and occulusion issues in recurrently refined RNN-like methods. By leveraging the Mamba’s powerful autoregressive capability and linear complexity, it offers significant potential to enhance accuracy and speed simultaneously through ablation studies.
2 Related Work
2.1 Optical Flow Estimation
@@
Optical flow estimation was traditionally formulated as an energy minimization task [1, 12, 38], with early variational methods like Horn and Schunck [12] relying on brightness constancy and spatial smoothness assumptions. Subsequent optimization-based approaches improved robustness through regularization [2].
@@
In the era of deep learning, following the introduction of FlowNet [5] as the first CNN-based optical flow method, numerous subsequent algorithms such as PWC-Net [44] and SpyNet [40], adopted a coarse-to-fine strategy for optical flow estimation. However, these approaches struggled with fast motions and small displacements due to inaccurate coarse-level guidance. RAFT [45] addressed this via iterative full-field refinement, inspiring variants such as GMA [20], SeparableFlow [62],and MS-RAFT [19], which introduced modules to better handle occlusion and motion ambiguity. Despite their accuracy, these CNN-based methods require heavy iterative computation and struggle with large displacements and complex motion, limiting their efficiency on resource-constrained devices.
@@
Transformers have emerged as powerful alternatives due to their ability to model global dependencies via self-attention. Models like FlowFormer [15], CRAFT [43], and GAFlow [31] leveraged Transformer architectures to enhance feature expressiveness and global matching. GMFlow [54] and TransFlow [30] further improved efficiency by reformulating optical flow as a one-shot global matching problem. However, the quadratic complexity of attention mechanisms results in high computational and memory costs, limiting deployment on lightweight systems.
2.2 State Space Models
@@
State space models [11], which originated from control theory, have been effectively integrated with deep learning to capture long-range dependencies. Mamba [8] introduced an input adaptation mechanism to enhance the state space model, providing faster inference speed, higher throughput and overall indicators compared to the Transformers of the same size, especially in natural language. Related work on Mamba [50] has shown that it has comparable performance to the Transformer-based methods while significantly reducing computational complexity. This superiority led to further exploration like Vision Mamba [68] and Vmamba [29], which applied SSMs to visual tasks such as image classification [29, 68], video understanding [25, 47], biomedical image segmentation [34, 53] and others [14, 17, 37, 39, 61]. However, their applicability to optical flow estimation has remained unexplored until recently. Lin et al. [26] pioneered the use of Mamba in the scene flow estimation task, proposing an SSM-based iterative update module to replace the GRU module as a decoder. Nevertheless, it was specifically designed for scene flow with point cloud feature and is not suitable for the generally used backbone optical flow methods.
@@
Inspired by the Mamba’s strength of long sequence modeling and autoregression with linear complexity, we propose a Mamba-centric architecture for optical flow estimation. Our goal is to design an end-to-end solution that utilizes Mamba for feature enhancement and flow optimization. During the feature enhancement phase, PolyMamba enhances features from pairs of images by sequentially employing Self-Mamba, Cross-Mamba, and MLP to capture global information and mutual dependencies. At the flow refinement phase, PulseMamba iteratively optimize the fina optical flow by leveraging a motion encoder and a SSM. These Mamba-centric designs guarantee the accuracy of our method while maintaining its high speed in training and inference, providing valuable insights for resource-constrained devices.
@@ 3.1 Preliminaries
3.2 Overall Architecture
Optical flow can be intuitively understood as a task aimed at finding pixel-wise correspondences between consecutive images. As shown in Fig. 2, our proposed framework, MambaFlow, is centered around two novel modules: the feature enhancement module PolyMamba and the flow refinement module PulseMamba.
@@ 3.3 Self-Mamba
[…]
@@ 3.4 Cross-Mamba
$\boldsymbol{x}_{\mathrm{conv}}=\texttt{SiLU}\left(\texttt{DWConv}(\boldsymbol{\mathit{F}}_{1}^{\text{self}})\right),$ (11)
@@
$\boldsymbol{x}_{\mathrm{mod}}=\texttt{Proj}(\boldsymbol{\mathit{F}}_{2}^{\text{self}}).$ (12)
@@
$[\,\boldsymbol{\Delta},\;\mathbf{B},\;\mathbf{C}\,]=\texttt{Linear}([\boldsymbol{x}_{\mathrm{conv}};\;\boldsymbol{x}_{\mathrm{mod}}]),$ (13)
@@
$\boldsymbol{\mathit{F}}^{\text{cross}}=\texttt{OutProj}(\boldsymbol{y}).$ (14)
@@
[…]
3.5 Global Matching
@@
3.6 PulseMamba
To effectively address the challenges of occlusions and boundary conditions in optical flow estimation, we propose a novel module named PulseMamba. PulseMamba integrates cost volume information and motion features, achieving dynamic and context-aware optical flow propagation through iterative refinement.
Given the initial optical flow $\mathbf{V}_{\text{initial}}$, the source image features $\boldsymbol{F}_{q}$, and the 4D cost volume (denoted as $\mathbf{C}$), PulseMamba proceeds as follows:
First, motion features are extracted by encoding the initial flow and the 4D cost volume:
$\boldsymbol{M}=\text{MotionEncoder}(\mathbf{V}_{\text{initial}},\mathbf{C}),$ (18)
PulseMamba employs multiple stacked Mamba layers for iterative feature refinement. At each iteration $i$, the input is formed by concatenating the image features, motion features, and current flow prediction, and the flow is updated as:
$\mathbf{V}^{(i)}=\text{Mamba}^{(i)}\left([\boldsymbol{F}_{q};\;\boldsymbol{M};\;\mathbf{V}^{(i-1)}]\right),$ (19)
where $\mathbf{V}^{(0)}=\mathbf{V}_{\text{initial}}$.
The output flow from each iteration becomes the input for subsequent refinement. After $N$ iterations, the final refined optical flow is obtained:
@@ 3.6 PulseMamba
Through this iterative and context-aware refinement, PulseMamba significantly improves the accuracy and robustness of optical flow estimation, particularly under complex conditions involving occluded regions and ambiguous boundaries.
Table 1: Quantitative results on Sintel and KITTI. MambaFlow achieves at least 18% faster inference on the KITTI (test) dataset compared to other state-of-the-art methods with competitive performance. All compared methods were evaluated under the same training configuration, and a unified "C+T+S+K+H" training strategy was adopted to ensure the fairness of the results. For performance measurements, we used the normalized setting of batch size 1 and input resolution 540 × 960. In addition, all inference times are measured on the RTX 3090. The significant performance improvement highlights MambaFlow’s advantages in speed and efficiency. It is worth noting that while maintaining competitive accuracy on the Sintel and KITTI datasets, we have successfully achieved a significant improvement in speed, further verifying the superiority of MambaFlow in the speed-accuracy balance.
Training Data Method Sintel KITTI Inference Cost
[…]
C+T+S+K+H RAFT [45] 1.61 2.86 5.10 5.3M 140.7ms
SeparableFlow [62] 1.50 2.67 4.64 6.0M 229.1ms
GMA [20] 1.39 2.47 5.15 5.9M 183.3ms
DIP [66] 1.44 2.83 4.21 5.4M 498.9ms
GMFlowNet [64] 1.39 2.65 4.79 9.3M 244.3ms
CRAFT [43] 1.45 2.42 7.79 6.3M 483.4ms
MS-RAFT [18] 1.37 2.67 4.88 13.5M 627.3ms
MatchFlow [4] 1.33 2.64 4.72 14.8M 283.8ms
MS-RAFT+ [19] 1.23 2.68 4.15 16.0M 1108.2ms
GMFlow [54] 1.74 2.90 9.32 4.7M 138.5ms
MambaFlow (ours) 1.43 2.45 8.06 11.2M 113.5ms
4 Experiments
@@
Following the previous methodology, we first pre-train on the FlyingChairs (Chairs)[5] and FlyingThings3D (Things)[35] datasets. For fine-tuning, we adopt a mixed training set consisting of KITTI [36], HD1K [22], FlyingThings3D, and Sintel [3] training sets. Finally, we further fine-tune the model on the KITTI training set and report the performance on the respective online benchmarks.
@@
Our model is constructed using the PyTorch library and trained on 8 NVIDIA GeForce RTX 3090 GPUs. Following the training strategy from previous works [54, 45], we use the AdamW optimizer. The model is first pre-trained on the Chairs dataset for 100k iterations with a batch size of 16 and a learning rate of $4\times 10^{-4}$. Then, it is fine-tuned on the Things dataset for 200k iterations with a batch size of 8 and a learning rate of $2\times 10^{-4}$. During this process, we select the best-performing model on the Things dataset and fine-tune it for an additional 800k iterations to validate the ablation study. Subsequently, the model fine-tuned for 200k iterations on the Things dataset is used as the initial model and further trained on the Sintel dataset, keeping the same hyperparameters as those used for the Things dataset, including a batch size of 8 and a learning rate of $2\times 10^{-4}$. Finally, we continue training on the KITTI-2015 training set for 200k iterations with the batch size and learning rate unchanged.
Figure 3: Visual results on Sintel test set.
Table 2: PolyMamba components. Ablation study on individual components, revealing that Cross-Mamba contributes the most.
setup Things (val) Sintel (train) Param (M)
@@
full 2.91 - 1.23 - 2.76 - 11.2 -
w/o self 3.11 $\uparrow$6.91.27 $\uparrow$3.32.82 $\uparrow$2.28.9 $\downarrow$24.6%
w/o cross 5.01 $\uparrow$72.22.08 $\uparrow$69.14.18 $\uparrow$51.45.4 $\downarrow$54.2%
w/o MLP 3.22 $\uparrow$10.71.33 $\uparrow$8.12.96 $\uparrow$7.28.2 $\downarrow$30.5%
w/o pos 3.00 $\uparrow$3.11.26 $\uparrow$2.42.82 $\uparrow$2.211.2 -
Table 3: Number of PolyMamba blocks. Using 8 blocks of Mamba strikes an optimal balance between accuracy and parameter efficiency.

### v3

@@ Abstract
Recently, the Mamba architecture has demonstrated significant successes in various computer vision tasks, such as classification and segmentation. However, its application to optical flow estimation remains unexplored. In this paper, we introduce MambaFlow, a novel framework designed to leverage the high accuracy and efficiency of the Mamba architecture for capturing locally correlated features while preserving global information in end-to-end optical flow estimation. To our knowledge, MambaFlow is the first architecture centered around the Mamba design tailored specifically for optical flow estimation. It comprises two key components: (1) PolyMamba, which enhances feature representation through a dual-Mamba architecture, incorporating a Self-Mamba module for intra-token modeling and a Cross-Mamba module for inter-modality interaction, enabling both deep contextualization and effective feature fusion; and (2) PulseMamba, which leverages an Attention Guidance Aggregator (AGA) to adaptively integrate features with dynamically learned weights in contrast to naive concatenation, and then employs the intrinsic recurrent mechanism of Mamba to perform autoregressive flow decoding, facilitating efficient flow information dissemination. Extensive experiments demonstrate that MambaFlow achieves remarkable results comparable to mainstream methods on benchmark datasets. Compared to SEA-RAFT, MambaFlow attains higher accuracy on the Sintel benchmark, demonstrating stronger potential for real-world deployment on resource-constrained devices. The source code will be made publicly available upon acceptance of the paper.
Introduction
@@
Optical flow estimation, a fundamental problem in computer vision, aims to compute the motion vectors of pixels between consecutive frames in a video sequence. It holds significant and widespread applications in areas such as action recognition (Zhao et al. 2020; Xu et al. 2019), video interpolation (Huang et al. 2022b; Wu et al. 2022; Liu, Liu, and Lin 2020), autonomous driving (Yi et al. 2023), etc.
@@
Figure 1: End-point-error on Sintel (clean) vs. inference time (ms) and model params (M). Compared to state-of-the-art models, our proposed model achieves comparable accuracy while significantly reducing inference time, demonstrating an optimal balance between computational efficiency and performance.
Architectures based on Convolutional Neural Networks (CNNs) have achieved pioneering progress in the field of optical flow. FlowNet (Dosovitskiy et al. 2015), as the first end-to-end CNNs-based flow network, revealed the immense potential of this learning paradigm. However, it faced limitations in handling large displacements and complex motions. Subsequently, PWC-Net (Sun et al. 2018) improved estimation accuracy in complex motion scenarios by incorporating pyramid feature extraction and a cost-volume construction, which effectively captured multi-scale local features. Nevertheless, it still encountered challenges in large displacement scenarios. To address these issues, RAFT (Teed and Deng 2020) introduced a recurrent iterative refinement mechanism, which significantly enhanced the accuracy of optical flow estimation. However, such innovation brought a new challenge, which its optimization process relies on numerous iterations, leading to increased parameter counts and longer inference time. Many subsequent variants of the RAFT algorithm have attempted upgrates (Jahedi et al. 2024b; Wang, Lipson, and Deng 2025; Zheng et al. 2022; Sun et al. 2022; Dong, Cao, and Fu 2023; Jahedi et al. 2024a), such as MS-RAFT and SEA-RAFT, but their ability to handle large displacements and complex motions still requires further improvement. To tackle the aforementioned issues, the Transformer architecture, known for its powerful global modeling capabilities, has been introduced into optical flow estimation tasks (Huang et al. 2022a; Shi et al. 2023; Lu et al. 2023; Sui et al. 2022). Most of them leverage attention mechanism to enhance feature representation, which capture long-range dependencies to significantly improve the accuracy. Notably, unlike optical flow algorithms based on RNNs for flow decoding (Teed and Deng 2020; Jiang et al. 2021), GMFlow (Xu et al. 2022) utilizes the attention mechanism to iteratively decode the final optical flow, resulting in faster flow refinement. However, they encounter the quadratic computational complexity of Transformers both during the training and inference stages (Xu et al. 2022; Sui et al. 2022; Huang et al. 2022a; Shi et al. 2023; Lu et al. 2023), which severely limits their efficiency in practical applications.
@@
Recently, Mamba (Gu and Dao 2023; Gu et al. 2021) has been widely used in various computer vision tasks, such as image classification (Ge et al. 2024; Yang, Wang, and Chen 2024; Yang et al. 2024),image segmentation (Liu et al. 2024a; Zhou et al. 2024; Du, Wang, and Chen 2024), object detection (Ma et al. 2024; Ma, Zhang, and Shi 2024; Verma et al. 2024), image generation (Hu et al. 2024; Lambrechts et al. 2024), video understanding (Jin et al. 2024; Li et al. 2024; Xiao et al. 2024), etc. Those remarkable achievements have demonstrated that Mamba can effectively address the challenges posed by CNNs and Transformers, including the local perception limitations of CNNs and the quadratic computational complexity of Transformers (Xu et al. 2024; Zhang et al. 2024). It benefits from a selective state space mechanism, which can efficiently capture global information and handle complex spatiotemporal dependencies with linear complexity (Zhang et al. 2024). Specifically, Mamba has been conceptually summarized as being ideally suited for tasks involving long sequences and autoregression (Yu and Wang 2024) due to its inherent RNN mechanism within the SSM (Gu et al. 2021; Gu, Goel, and Ré 2021; Gu and Dao 2023).
@@
Inspired by Mamba’s capabilities for long sequence modeling and autoregression with linear complexity, we propose a novel optical flow estimation framework, MambaFlow, designed to address the challenges in CNNs or Transformer-based methods. As demonstrated in Figure 1, our method significantly reduces inference time while maintaining remarkblely high accuracy.The framework introduces two core components: (1) the PolyMamba module, which utilizes Self-Mamba, Cross-Mamba, and Multi-Layer Perceptron (MLP) to globally optimize feature representations with lightweight computation, and (2) the PulseMamba module, which first employs an Attention Guidance Aggregator (AGA) to adaptively integrate motion features using dynamically learned weights—instead of naive concatenation—and then leverages Mamba’s autoregressive capability to iteratively decode and refine optical flow, effectively addressing challenges such as large displacement and occluded regions.
@@
Specifically, we begins with a CNN backbone to extract robust features from the input image pairs similar to traditional practice (Wang, Lipson, and Deng 2025; Xu et al. 2022; Jiang et al. 2021). Subsequently, we leverage the suggested PolyMamba to enhance the extracted features, allowing for the capture of complex global information and long-range dependencies. Building on the enhanced features, we further utilize a global matching mechanism (Xu et al. 2022; Zhao et al. 2022) to compute similarities for initial flow. At last, we employ PulseMamba to recurrently refine and yield the final precise flow (Xu et al. 2022). Our algorithm uses Mamba as the core components in both the encoding and decoding stage, making it an end-to-end optical flow estimation algorithm centered around Mamba. By leveraging Mamba’s capabilities in global modeling and autoregressive properties, as well as its computationally efficient characteristics, the proposed approach achieves a good trade-off between accuracy and performance.
@@
  • We design the PulseMamba, which integrates an Attention Guidance Aggregator for adaptive feature fusion and a Mamba-based autoregressive decoder, addressing the computational overload, large displacement and occulusion issues in recurrently refined RNN-like methods. By leveraging the Mamba’s powerful autoregressive capability and linear complexity, it offers significant potential to enhance accuracy and speed simultaneously through ablation studies.
Related Work
Optical Flow Estimation
@@
Optical flow estimation was traditionally formulated as an energy minimization task (Brox et al. 2004; Horn and Schunck 1981; Papenberg et al. 2006), with early variational methods like Horn and Schunck (Horn and Schunck 1981) relying on brightness constancy and spatial smoothness assumptions. Subsequent optimization-based approaches improved robustness through regularization (Bruhn, Weickert, and Schnörr 2005).
@@
In the era of deep learning, following the introduction of FlowNet (Dosovitskiy et al. 2015) as the first CNN-based optical flow method, numerous subsequent algorithms such as PWC-Net (Sun et al. 2018) and SpyNet (Ranjan and Black 2017), adopted a coarse-to-fine strategy for optical flow estimation. However, these approaches struggled with fast motions and small displacements due to inaccurate coarse-level guidance. RAFT (Teed and Deng 2020) addressed this via iterative full-field refinement, inspiring variants such as GMA (Jiang et al. 2021), SeparableFlow (Zhang et al. 2021), and MS-RAFT (Jahedi et al. 2024b), which introduced modules to better handle occlusion and motion ambiguity. Despite their accuracy, these CNN-based methods require heavy iterative computation and struggle with large displacements and complex motion, limiting their efficiency on resource-constrained devices.
@@
Transformers have emerged as powerful alternatives due to their ability to model global dependencies via self-attention. Models like FlowFormer (Huang et al. 2022a), CRAFT (Sui et al. 2022), and GAFlow (Luo et al. 2023) leveraged Transformer architectures to enhance feature expressiveness and global matching. GMFlow (Xu et al. 2022) and TransFlow (Lu et al. 2023) further improved efficiency by reformulating optical flow as a one-shot global matching problem. However, the quadratic complexity of attention mechanisms results in high computational and memory costs, limiting deployment on lightweight systems.
State Space Models
@@
State space models (Gu, Goel, and Ré 2022), which originated from control theory, have been effectively integrated with deep learning to capture long-range dependencies. Mamba (Gu and Dao 2023) introduced an input adaptation mechanism to enhance the state space model, providing faster inference speed, higher throughput and overall indicators compared to the Transformers of the same size, especially in natural language. Related work on Mamba (Wu et al. 2025) has shown that it has comparable performance to the Transformer-based methods while significantly reducing computational complexity. This superiority led to further exploration like Vision Mamba (Zhu et al. 2024) and Vmamba (Liu et al. 2024b), which applied SSMs to visual tasks such as image classification (Liu et al. 2024b; Zhu et al. 2024), video understanding (Li et al. 2025; Wang et al. 2023), biomedical image segmentation (Ma, Li, and Wang 2024; Xing et al. 2024) and others (Hu et al. 2025; Islam et al. 2023; Nguyen et al. 2022; Qin et al. 2024; Zha et al. 2024). However, their applicability to optical flow estimation has remained unexplored until recently. Lin et al. (Lin et al. 2024) pioneered the use of Mamba in the scene flow estimation task, proposing an SSM-based iterative update module to replace the GRU module as a decoder. Nevertheless, it was specifically designed for scene flow with point cloud feature and is not suitable for the generally used backbone optical flow methods.
@@
Inspired by the Mamba’s strength of long sequence modeling and autoregression with linear complexity, we propose a Mamba-centric architecture for optical flow estimation. Our goal is to design an end-to-end solution that utilizes Mamba for feature enhancement and flow optimization. These Mamba-centric designs guarantee the accuracy of our method while maintaining its high speed in training and inference, providing valuable insights for resource-constrained devices.
@@
Overall Architecture
Optical flow can be intuitively understood as a task aimed at finding pixel-wise correspondences between consecutive images. As shown in Figure 2, our proposed framework, MambaFlow, is centered around two novel modules: the feature enhancement module PolyMamba and the flow refinement module PulseMamba.
@@
$\boldsymbol{\mathit{F}}_{i}^{\text{self}}$ $=\text{Self-Mamba}(\boldsymbol{\mathit{F}}_{i})$ (10)
$=\text{Self-Forward}(\boldsymbol{\mathit{F}}_{i})+\text{Self-Backward}(\boldsymbol{\mathit{F}}_{i})$
$\quad i\in\{1,2\}.$
@@
$\boldsymbol{x}_{\mathrm{conv}}=\operatorname{SiLU}\left(\operatorname{DWConv}(\boldsymbol{\mathit{F}}_{1}^{\text{self}})\right),$ (11)
@@
$\boldsymbol{x}_{\mathrm{mod}}=\operatorname{Proj}(\boldsymbol{\mathit{F}}_{2}^{\text{self}}).$ (12)
@@
$\boldsymbol{\Delta},\;\mathbf{B},\;\mathbf{C}\,]=\operatorname{Linear}([\boldsymbol{x}_{\mathrm{conv}};\;\boldsymbol{x}_{\mathrm{mod}}]),$ (13)
@@
$\boldsymbol{\mathit{F}}^{\text{cross}}=\operatorname{OutProj}(\boldsymbol{y}).$ (14)
@@
$\text{Cross-Mamba}\left(\boldsymbol{\mathit{F}}_{1}^{\text{self}},\;\boldsymbol{\mathit{F}}_{2}^{\text{self}}\right)$ (15)
$=\;\text{Forward}(\,\cdot\,)+\text{Backward}(\,\cdot\,).$
Global Matching
@@
Extra Data Method Sintel KITTI Inference Cost
[…]
RAFT ECCV’20 1.61 2.86 5.10 5.3M 140.7ms
SeparableFlow ICCV’21 1.50 2.67 4.64 6.0M 229.1ms
GMA ICCV’21 1.39 2.47 5.15 5.9M 183.3ms
DIP CVPR’22 1.44 2.83 4.21 5.4M 498.9ms
GMFlowNet CVPR’22 1.39 2.65 4.79 9.3M 244.3ms
CRAFT CVPR’22 1.45 2.42 4.79 6.3M 483.4ms
GMFlow CVPR’22 1.74 2.90 9.32 4.7M 138.5ms
SKFlow NeurIPS’22 1.28 2.23 4.85 6.3M 331.9ms
MatchFlow CVPR’23 1.33 2.64 4.72 14.8M 283.8ms
MS-RAFT+ IJCV’24 1.23 2.68 4.15 16.0M 1108.2ms
MambaFlow (ours) 1.20 2.62 7.34 20.5M 116.5ms
CroCo-Pretrain CroCoFlow ICCV’23 1.09 2.44 3.64 437.4M 6422.0ms
DDVM-Pretrain DDVM NeurIPS’23 1.75 2.48 3.26
YouTube-VOS FlowFormer++ CVPR’23 1.07 1.94 4.52 16.1M 373.4ms
VIPER CCMR+ WACV’24 1.07 2.10 3.86 11.5M OOM
TartanAir SEA-RAFT(M) ECCV’24 1.44 2.86 4.64 18.8M 71.0ms
TartanAir SEA-RAFT(L) ECCV’24 1.31 2.60 4.30 18.8M 108.0ms
Table 1: Quantitative results on Sintel and KITTI benchmarks. All methods are trained on the unified C+T+S+K+H datasets as the baseline. The ”Extra Data” column indicates additional pretraining datasets. Evaluation is conducted with batch size 1, input resolution $540\times 960$, and an RTX 3090 GPU.
Figure 3: Visual results on Sintel test set.
PulseMamba
To effectively address occlusions and ambiguous boundary regions in optical flow estimation, we propose PulseMamba, an autoregressive refinement module featuring an Attention Guidance Aggregator (AGA) for adaptive feature integration. While conventional feature concatenation equally weights all spatial information, it often neglects dynamically relevant regions. To address this limitation, AGA adaptively assigns spatial attention weights to each feature, enabling context-aware flow estimation.
Given initial optical flow $\mathbf{V}_{\text{initial}}$, source image features $\boldsymbol{F}_{q}$, and a 4D correlation cost volume $\mathbf{C}$, PulseMamba first extracts motion features:
$\boldsymbol{M}=\text{MotionEncoder}(\mathbf{V}_{\text{initial}},\mathbf{C}).$ (18)
AGA adaptively integrates motion features $\boldsymbol{M}$, context features $\boldsymbol{F}_{q}$, and hidden-state features $\boldsymbol{h}$ into a contextually informative representation. Specifically, AGA first aligns these input features to a unified embedding dimension through individual convolutional projections.
Then, concatenated aligned features are processed via convolutional and GELU layers to produce spatial attention maps $\boldsymbol{A}$:
$\boldsymbol{M};\,\boldsymbol{F}_{q};\,\boldsymbol{h}]\xrightarrow{\mathrm{Conv}_{1}}\xrightarrow{\mathrm{GELU}}\xrightarrow{\mathrm{Conv}_{2}}\xrightarrow{\mathrm{Softmax}}\boldsymbol{A.}$ (19)
Finally, the aggregated feature $\boldsymbol{x}_{\text{AGA}}$ is obtained by a weighted sum:
$\boldsymbol{x}_{\text{AGA}}=\sum_{f\in\{M,F_{q},h\}}\boldsymbol{A}_{f}\odot\boldsymbol{f},$ (20)
where $\boldsymbol{A}_{f}$ represents spatial attention weights corresponding to each feature type.
The aggregated feature $\boldsymbol{x}_{\text{AGA}}$ is then iteratively refined by PulseMamba using a Mamba layer to update the hidden-state representation:
$\boldsymbol{h}^{(i)}=\text{Mamba}(\boldsymbol{x}_{\text{AGA}}^{(i-1)}).$ (21)
Subsequently, the flow increment $\Delta\mathbf{V}^{(i)}$ is predicted, and the optical flow is iteratively updated:
$\Delta\mathbf{V}^{(i)}$ $=\text{FlowHead}(\boldsymbol{h}^{(i)})$ (22)
$\mathbf{V}^{(i)}$ $=\mathbf{V}^{(i-1)}+\Delta\mathbf{V}^{(i)}.$
After NN iterations, PulseMamba produces the final refined optical flow:
@@
Through dynamic attention-guided feature integration and autoregressive refinement, PulseMamba effectively enhances the robustness and precision of optical flow estimation, especially under challenging scenarios.
Experiments
@@
Following the previous methodology, we first pre-train on the FlyingChairs (Chairs)(Dosovitskiy et al. 2015) and FlyingThings3D (Things)(Mayer et al. 2016) datasets. For fine-tuning, we adopt a mixed training set consisting of KITTI (Menze and Geiger 2015), HD1K (Kondermann et al. 2016), FlyingThings3D, and Sintel (Butler et al. 2012) training sets. Finally, we further fine-tune the model on the KITTI training set and report the performance on the respective online benchmarks.
@@
Our model is constructed using the PyTorch library and trained on 8 NVIDIA GeForce RTX 3090 GPUs. Following the training strategy from previous works (Xu et al. 2022; Teed and Deng 2020), we use the AdamW optimizer. The model is first pre-trained on the Chairs dataset for 100k iterations with a batch size of 16 and a learning rate of $4\times 10^{-4}$. Then, it is fine-tuned on the Things dataset for 200k iterations with a batch size of 8 and a learning rate of $2\times 10^{-4}$. During this process, we select the best-performing model on the Things dataset and fine-tune it for an additional 800k iterations to validate the ablation study. Subsequently, the model fine-tuned for 200k iterations on the Things dataset is used as the initial model and further trained on the Sintel dataset, keeping the same hyperparameters as those used for the Things dataset, including a batch size of 8 and a learning rate of $2\times 10^{-4}$. Finally, we continue training on the KITTI-2015 training set for 200k iterations with the batch size and learning rate unchanged.
Comparison with State-of-the-Arts
@@
Among the methods that utilize extra pretraining data, SEA-RAFT is recognized as one of the fastest baselines. Our model achieves a similar inference speed to SEA-RAFT(L). Despite the use of additional TartanAir data in SEA-RAFT(L), MambaFlow achieves a lower EPE on Sintel Clean, with an 8.4% reduction compared to SEA-RAFT(L). For other methods that utilize large-scale extra pretraining data, such as FlowFormer++ (YouTube-VOS(Xu et al. 2018)), CroCoFlow(Weinzaepfel et al. 2023) (CroCo-Pretrain), and others, MambaFlow achieves comparable accuracy while maintaining significantly faster inference.
Figure 3 presents a qualitative analysis of optical flow estimation, comparing MambaFlow with other state-of-the-art approaches on the Sintel test set.
Overall, these results demonstrate that MambaFlow achieves a favorable trade-off between accuracy and efficiency, and delivers performance comparable to state-of-the-art methods, regardless of whether additional pretraining data is utilized.
Setup Things (val) Sintel (train) Param
clean clean final (M)
full 2.22 1.15 3.05 20.5
w/o self 2.40 1.24 3.12 18.2
w/o cross 3.84 2.00 4.61 14.7
w/o MLP 2.46 1.25 3.16 17.5
w/o pos 2.29 1.18 3.11 20.5
Table 2: Ablations on the Components of PolyMamba